Driving down a mountain, George finds that he has descended 1800 ft in elevation by the time he is 3.25 miles horizontally away from the top of the mountain. Find the slope of his descent to the nearest hundredth.
Answer by checkley77(12844)

(0,0) & (3.25*5,280,1,800)
(0,0) & (17,160,1,800)
slope=(y2-y1)/(x2-x1)
slope=(1,800-0)/(17,160-0)
slope=1,800/17,160
slope=.1049 ans.